X and Y are dependent random variable with E(X)=1, E(Y)=6, Var(X)=2, and Var(Y)=2 find the following:

a. E(X-6)

b. E(XY)

c. E(X+Y)


1
Expert's answer
2022-01-17T12:00:04-0500

Given that "E(X)=1, E(Y)=6, Var(X)=2, and\space Var(Y)=2" then, 

"a)"

"E(X-6)=E(X)-E(6)=1-6=-5," since the expectation of a constant is the same constant.

"b)"

To find "E(XY)" , we shall derive it from the covariance of the random variables "X" and "Y"

Now,

"Cov(x,y)=E(xy)-(E(x)\times E(y))"

We make "E(xy)" subject of the formula.

So,

"E(xy)=Cov(x,y)+(E(x)\times E(y))"

Given that "E(X)=1\space and \space E(Y)=6" then,

"E(xy)=Cov(x,y)+(1\times 6)\\
E(xy)=Cov(x,y)+6..........(1)"

Given the value of "Cov(x,y)", we can find the value of "E(xy)" using equation (1) above.


"c)"

"E(X+Y)=E(X)+E(Y)=1+6=7," this is because, expectation of the sum of the two variables is equal to the sum of the mathematical expectation of "X" and the mathematical expectation of "Y".